# Verifiable mission planning for space operations

### v2

@@ Abstract
The growing complexity of space missions necessitates planning methods that maximize mission performance while rigorously enforcing safety. We develop a probabilistic approach based on a finite-horizon Markov decision process to optimize spacecraft operations planning with safety guarantees. In the model, states capture essential mission parameters, and actions represent the operational adjustments needed to meet mission objectives. By directly incorporating uncertainties from environmental conditions and spacecraft dynamics, an optimal sequence of actions is computed that maximizes expected rewards and strictly enforces safety constraints. Numerical experiments on the GRACE-FO mission demonstrate robust performance under uncertainties while providing probabilistic safety guarantees, offering a reliable solution for autonomous spacecraft operations.
@@ 1 Introduction
Spacecraft mission planning is increasingly challenged by the complex uncertainties inherent in space operations. Modern missions must perform high-precision observations while contending with unpredictable space weather, limited fuel reserves, and intricate orbital dynamics. Traditionally, mission planning relied on heuristic methods studying static scenarios to account for these uncertainties [1]. However, such approaches often fail to capture the full probabilistic nature of environmental variability and spacecraft dynamics, leading to suboptimal performance or potential safety violations.
Markov decision processes (MDP) provide a promising alternative. MDP is a mathematical framework for modeling planning and acting in stochastic environments with nondeterministic action selection[2]. An MDP models the system as a collection of states, actions, and probabilistic transitions, with each action yielding a corresponding reward. In the context of spacecraft mission planning, the system’s states encapsulate key operational parameters—such as orbital altitude and fuel reserves—while the actions represent possible maneuvers or operational adjustments. By explicitly incorporating uncertainties, such as fluctuations in space weather and variations in orbital dynamics, MDPs enable the derivation of optimal policies that balance performance objectives with verifiable safety constraints.
The GRACE-FO mission[3] serves as a motivating scenario. Originally designed to measure Earth’s gravity field with high precision by flying two satellites in tandem, the mission soon faced critical setbacks when one accelerometer failed and a thruster leak emerged. Despite these challenges, GRACE-FO must maintain its primary objective by reducing atmospheric drag to preserve data quality and avoid resonant orbits, even though it operates on limited fuel and lacks provisions for altitude-raising maneuvers. Save et al. [1] introduced a traditional orbit management strategy for GRACE-FO that considers various solar activity scenarios; however, their method does not fully capture solar flux variability, sometimes resulting in premature reentry under worst-case conditions.
We propose to address these challenges by applying a finite-horizon MDP to space mission planning, which explicitly incorporates uncertainty into its state transitions. This framework enables adaptive policy synthesis that optimally schedules maneuvers based on probabilistic models of environmental conditions and spacecraft dynamics. In contrast to traditional scenario-based methods, the MDP-based approach continuously refines its decision-making process in response to real-time variations, thereby enhancing performance while rigorously enforcing safety constraints.
@@
Prior studies on mission planning under uncertainty have explored various methods. Early scenario-based analyses, as presented by Save et al.[1], relied on fixed worst-case, nominal, and best-case evaluations. However, this approach often results in overly conservative planning strategies. Recent work by Eddy et al.[4] applied MDPs to integrate uncertainty into the planning process, enabling dynamic balancing of multiple operational objectives. Other studies have taken alternative routes. Isaji and Ho[5] developed an integrated approach that combines stochastic programming with decomposition-based optimization to address both mission planning and spacecraft design under uncertainty. Vasile [6] leveraged evidence theory along with multi-agent collaborative search to generate robust mission designs by quantifying uncertain design parameters. Additionally, as explored by Ahmadi et al. [7], risk-averse planning techniques incorporate dynamic coherent risk measures to synthesize policies that explicitly mitigate adverse outcomes.
We present a general framework for probabilistic mission planning and safety verification for spacecraft operations. First, a finite-horizon MDP framework is developed that directly embeds uncertainties into state transitions, enabling adaptive and real-time decision making. Second, we present how to verify the safety of the obtained strategy regarding some operational constraints. We then describe how to construct this MDP-based solution for the case of GRACE-FO. Finally, we show that it not only improves fuel efficiency and data quality but also provides measurable safety guarantees. These contributions extend traditional scenario-based methods [1] and build on dynamic MDP approaches [4], offering a comprehensive and computationally efficient solution for robust spacecraft mission planning under uncertainty with safety guarantees.
@@ 2 Probabilistic Policy Synthesis and Safety Verification Using Markov Decision Processes
A Markov decision process (MDP) is a common paradigm for modeling planning and acting in stochastic environments with nondeterministic action selection [2]. For mission planning, we focus on finite-horizon MDPs, wherein the decision-making process is performed over a finite number of time steps called a planning horizon. In the context of a spacecraft mission, the environment stochasticity could be due to, e.g., uncertainties in future space weather conditions, while actions could correspond to performing altitude raises. Similarly, the planning horizon could be the operational lifetime of the mission, while individual time steps are specific points in time at which actions are chosen, e.g., daily, weekly, or monthly.
Formally, a finite-horizon MDP $\mathfrak{M}$ is described by the tuple $\mathfrak{M}=\langle\mathcal{S},\mathcal{A},\mathcal{H},T,R,\mathbf{P}_{0}\rangle$, where $s\in\mathcal{S}$ is a finite set of states, $a\in\mathcal{A}$ is a finite set of actions, =$\mathcal{H}=\{0,1,2,\ldots,h,\ldots,H\}$ is a set of time indices with planning horizon $H<\infty$, $T:\mathcal{S}\times\mathcal{A}\times\mathcal{H}\times\mathcal{S}\rightarrow[0,1]$, is a transition function mapping state-action pairs at time step $h$ to a probability distribution over successor states, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{H}\rightarrow\mathbb{R}$ is a reward function mapping state-action pairs at time step $h$ to an associated reward, and $\mathbf{P}_{0}:\mathcal{S}\rightarrow[0,1]$ is an initial state distribution. Let $R(s,a,h)$ denote the reward obtained for selecting action $a$ while in state $s$ at time step $h$. Similarly, let $T(s^{\prime}|a,s,h)$ denote the probability of transitioning to state $s^{\prime}$ after taking action $a$ while in state $s$ at time step $h$. Note that since the transition function must define a valid probability distribution, we have $\sum_{s^{\prime}\in\mathcal{S}}T(s^{\prime}|a,s,h)=1$ for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times\mathcal{H}$.
@@
To resolve the nondeterminism in choosing an action, the decision-maker must synthesize a time-dependent policy $\pi\in\Pi$   where $\Pi$ is the set of time-dependent policies. Formally, $\pi$ is a sequence of decision rules $\{d_{1},d_{2},\ldots,d_{H}\}$, where each $d_{h}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ prescribes a probability distribution for action selection in each state at time step $h$. Notice that each decision rule is a function of only the current state of the decision-maker, rather than of its entire history of states. Such a policy is referred to as Markovian. Let $\pi(a|s,h)$ denote the probability that action $a$ is selected while in state $s$ at time step $h$ under policy $\pi$ If there exists an action $a\in\mathcal{A}$ for each $(s,h)\in\mathcal{S}\times\mathcal{H}$ such that $\pi(a|s,h)=1$ and $\pi(a^{\prime}|s,h)=0$ for all $a^{\prime}\in\mathcal{A}\setminus a$, we refer to $\pi$ as deterministic. Finally, for a time-dependent policy, let $\pi(s,h)$ denote the action selected deterministically while in state $s$ at time step $h$.
@@
$\pi^{*}=\mathrm{arg}\max_{\pi\in\Pi}\mathbb{E}\left[\sum\nolimits_{h=1}^{H}R(s,a,h)\right].$ (4)
$\pi^{*}$ is obtained by solving the recursive Bellman equations backwards-in-time according to
@@
where $V^{*}(s,h)$ is the optimal value function for state $s$ at time step $h$ and $\pi^{*}(s,h)$ is the corresponding maximizing value of $a\in\mathcal{A}$. We then evaluate the safety performance of the obtained policy $\pi$. To do this, we define $V^{\varphi}(s,h)=\mathrm{Pr}^{\pi}((s,h)\models\square\neg\mathcal{B}).$ Note that here $\pi$ is the policy obtained from solving the unconstrained problem (4). The safety value function $V^{\varphi}(s,h)$ is computed using the following formulation:
@@
Note that it is straightforward to generalize the constraint (13b) to the probability of multiple mission specifications. We refer the reader to [10], section 8 for further details.
@@ 3 Space Environment and Dynamics
Transitioning from the MDP formulation to a detailed representation of the operational environment is crucial for accurate policy synthesis. Although the MDP framework guides decision making under uncertainty, its effectiveness relies on a clear understanding of the forces affecting spacecraft motion. Orbital mechanics forms the foundation for modeling a spacecraft’s trajectory by accounting for gravitational forces, atmospheric drag, and other perturbations, which in turn define the deterministic states, actions, and transitions in our MDP. At the same time, space weather conditions directly impact the environment in which the spacecraft operates, yet their inherent variability makes them very difficult to predict accurately. These uncertainties are incorporated into the transition probabilities of the model.
@@ 3.1 Orbital Mechanics
First, we introduce key ideas from orbital mechanics to then present how we can build an MDP for space mission planning. Consider a spacecraft in an elliptical orbit around the Earth. Let $\mathbf{r}(t)=[\mathbf{p}(t)^{\top},\mathbf{v}(t)^{\top}]^{\top}\in\mathbb{R}^{6}$ denote the Cartesian state of a spacecraft in the inertial frame, where $\mathbf{p}(t)=[x(t),y(t),z(t)]^{\top}$ and $\mathbf{v}(t)=[\dot{x}(t),\dot{y}(t),\dot{z}(t)]^{\top}$ are the spacecraft position and velocity at time instant $t$. In what follows, we will omit the time index $t$ for notational clarity. Assuming a spherical Earth, the altitude $R_{\mathrm{alt}}$ of the spacecraft above Earth’s surface can subsequently be obtained from its position $\mathbf{p}$ according to
$R_{\mathrm{alt}}=||\mathbf{p}||_{2}-R_{e},$ (8)
where $R_{e}$ is the radius of the Earth.
The dynamics of the spacecraft are given by
$\ddot{\mathbf{r}}=-\mu\frac{\mathbf{r}}{||\mathbf{r}||_{2}^{3}}+F_{\mathrm{drag}}+F_{\mathrm{pert}},$ (9)
@@
where $C_{D}$ is the drag coefficient of the spacecraft, $A$ is the cross-sectional area of the spacecraft, $m$ is the mass of the spacecraft, $\rho_{A}$ is the altitude-dependent atmospheric density, in $\mathrm{kg}/\mathrm{km}^{3}$, $\mathbf{\textit{v}}_{a}$ is the velocity of the spacecraft relative to the Earth-fixed frame, given by
$\mathbb{\textit{v}}_{a}\triangleq\begin{bmatrix}\dot{x}+\dot{\theta}y\\ \dot{y}-\dot{\theta}x\\ \dot{z}\end{bmatrix},\quad v_{a}=||\mathbb{\textit{v}}_{a}||_{2},$
in which $\dot{\theta}$ is the rotation rate of the Earth. Here, we use the MSIS atmosphere model [11][12][13] to compute the value of $\rho_{A}$ as a function of the altitude, and the solar flux at 10.7 cm (F10.7) using NOAA prediction [14].
In tandem with the preceding Cartesian representation of the spacecraft dynamics, one can instead characterize its orbit in terms of the Keplerian orbital elements $\{a,e,i,\omega,\Omega,f\}$, where $a$ is the semimajor axis length of the elliptical orbit, $e\in[0,1)$ is the eccentricity of the orbit, $i$ is the inclination of the orbit, $\omega$ is the argument of periapse, $\Omega$ is the longitude of the ascending node, and $f$ is the true anomaly [15]. These orbital elements are visualized in Figure 2, where the angles are defined with respect to the Earth-centered inertial (ECI) reference frame.
@@ 3.2 Space Weather Conditions
The solar cycle governs the Sun’s activity levels and, in turn, the atmospheric environment encountered by the spacecraft. However, the duration and intensity of individual cycles vary; larger cycles tend to last longer than 11 years, while smaller cycles are typically shorter. Solar activity is commonly categorized into three levels: high, moderate, and low [1]. During high solar activity, increased ultraviolet and X-ray radiation heats and expands the Earth’s upper atmosphere, raising atmospheric density at orbital altitudes and increasing drag forces. This accelerates orbital decay, requiring more frequent station-keeping maneuvers and increasing fuel consumption. In moderate activity periods, atmospheric density variations are less extreme, allowing for more stable orbit maintenance. Conversely, low solar activity leads to reduced drag, minimizing fuel expenditure and extending mission lifetime.

### v3

@@ Abstract
Spacecraft must operate under environmental and actuator uncertainties while meeting strict safety requirements. Traditional approaches rely on scenario-based heuristics that fail to account for stochastic influences, leading to suboptimal or unsafe plans. We propose a finite-horizon, chance-constrained Markov decision process for mission planning, where states represent mission and vehicle parameters, actions correspond to operational adjustments, and temporal logic specifications encode operational constraints. We synthesize policies that optimize mission objectives while ensuring constraints are met with high probability. Applied to the GRACE-FO mission, the approach accounts for stochastic solar activity and uncertain thrust performance, yielding maneuver schedules that maximize scientific return and provably satisfy safety requirements. This work demonstrates how Markov decision processes can be applied to space missions, enabling autonomous operation with formal guarantees.
@@ 1 Introduction
Planning spacecraft operations is critical for sustaining mission objectives, as missions must adapt to variable conditions and limited resources. Traditional approaches rely on precomputed scenarios or heuristic scheduling rules that assume fixed environmental conditions[1]. These methods often fail to capture the stochastic nature of orbital decay and space weather variability, leading to inefficient fuel use or violations of operational constraints.
The GRACE-FO mission provides a motivating case study. Designed to map Earth’s gravity field through tandem satellite operations, it has faced two major challenges: malfunctioning thrusters limiting orbit control, and the failure of one satellite’s accelerometer. To mitigate the resulting degradation in measurement accuracy, the mission must operate at higher altitudes, where atmospheric drag is reduced. This requirement, combined with uncertain solar activity and fuel limitations, makes orbit planning particularly difficult. Previous orbit management strategies, such as those of Save et al. [1], rely on fixed solar cycle scenarios; however, they do not capture variability in solar flux and can result in premature reentry under adverse conditions.
We address these challenges by formulating spacecraft mission planning as a finite-horizon Markov decision process (MDP), a framework for sequential decision making under uncertainty [2]. For GRACE-FO, states capture mission-relevant quantities such as orbital altitude and fuel reserves, actions represent discrete maneuver options, and transition probabilities model the effects of uncertain factors like solar flux and thrust variability. The reward function encodes mission objectives, such as maintaining observation quality at higher altitudes and avoiding repeated groundtrack orbits. Solving the MDP yields maneuver policies that maximize mission return while enforcing constraints such as minimum altitude and maneuver spacing. Unlike traditional scenario-based methods, this approach directly incorporates uncertainty into the planning model, enabling adaptive maneuver scheduling with formal, probabilistic safety guarantees.
@@
Prior studies on mission planning under uncertainty have approached the problem from several directions. Scenario-based analyses, such as those of Save et al. [1], evaluate a fixed set of worst-case, nominal, and best-case conditions, but this often produces overly conservative strategies. Decision-theoretic approaches have sought to improve flexibility: Eddy et al. [3] applied MDPs to integrate uncertainty into satellite task planning, enabling trade-offs across multiple objectives. Other work has pursued optimization-based methods, such as Isaji and Ho [4], who combined stochastic programming with decomposition techniques to jointly address mission planning and spacecraft design. In parallel, robust design strategies have been explored, including evidence-theory formulations for quantifying uncertain parameters [5] and risk-averse planning techniques that incorporate coherent risk measures to mitigate adverse outcomes [6]. While these studies advance uncertainty-aware mission planning, they do not provide explicit probabilistic safety guarantees tailored to spacecraft operations, the focus of this work.
We develop a finite-horizon MDP method for probabilistic mission planning and safety verification in spacecraft operations. The model embeds environmental and system uncertainties directly into state transitions, enabling maneuver strategies that respect formal operational constraints. We apply this approach to the GRACE-FO mission, demonstrating how the resulting policies improve fuel efficiency and data quality while satisfying safety requirements such as minimum altitude. By extending scenario-based orbit management [1] and advancing decision-theoretic planning [3], our approach presents a proof of concept for applying formal safety guarantees to space mission planning, pointing toward certifiable autonomy in future missions.
@@ 2 Probabilistic Policy Synthesis and Safety Verification Using Markov Decision Processes
We model spacecraft decision-making under uncertainty as a finite-horizon Markov decision process (MDP). An MDP is a common paradigm for modeling planning and acting in stochastic environments with nondeterministic action selection [2]. For mission planning, we focus on finite-horizon MDPs, wherein the decision-making process is performed over a finite number of time steps called a planning horizon. In the context of a spacecraft mission, the environment stochasticity could be due to, e.g., uncertainties in future space weather conditions, while actions could correspond to performing altitude raises. Similarly, the planning horizon could be the operational lifetime of the mission, while individual time steps are specific points in time at which actions are chosen, e.g., daily, weekly, or monthly.
Formally, a finite-horizon MDP $\mathfrak{M}$ is described by the tuple $\mathfrak{M}=\langle\mathcal{S},\mathcal{A},\mathcal{H},T,R,\mathbf{P}_{0}\rangle$, where $s\in\mathcal{S}$ is a finite set of states, $a\in\mathcal{A}$ is a finite set of actions, $\mathcal{H}=\{0,1,2,\ldots,h,\ldots,H\}$ is a set of time indices with planning horizon $H<\infty$, $T:\mathcal{S}\times\mathcal{A}\times\mathcal{H}\times\mathcal{S}\rightarrow[0,1]$, is a transition function mapping state-action pairs at time step $h$ to a probability distribution over successor states, $R:\mathcal{S}\times\mathcal{A}\times\mathcal{H}\rightarrow\mathbb{R}$ is a reward function mapping state-action pairs at time step $h$ to an associated reward, and $\mathbf{P}_{0}:\mathcal{S}\rightarrow[0,1]$ is an initial state distribution. Let $R(s,a,h)$ denote the reward obtained for selecting action $a$ while in state $s$ at time step $h$. Similarly, let $T(s^{\prime}|a,s,h)$ denote the probability of transitioning to state $s^{\prime}$ after taking action $a$ while in state $s$ at time step $h$. Note that since the transition function must define a valid probability distribution, we have $\sum_{s^{\prime}\in\mathcal{S}}T(s^{\prime}|a,s,h)=1$ for all $(s,a,h)\in\mathcal{S}\times\mathcal{A}\times\mathcal{H}$.
@@
To resolve the nondeterminism in choosing an action, the decision-maker must synthesize a time-dependent policy $\pi\in\Pi$   where $\Pi$ is the set of time-dependent policies. Formally, $\pi$ is a sequence of decision rules $\{d_{1},d_{2},\ldots,d_{H}\}$, where each $d_{h}:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$ prescribes a probability distribution for action selection in each state at time step $h$. A policy is simply a rule that can specify, for example, which maneuver a spacecraft should execute given its current state and time. Notice that each decision rule is a function of only the current state of the decision-maker, rather than of its entire history of states. Such a policy is referred to as Markovian. Let $\pi(a|s,h)$ denote the probability that action $a$ is selected while in state $s$ at time step $h$ under policy $\pi$ If there exists an action $a\in\mathcal{A}$ for each $(s,h)\in\mathcal{S}\times\mathcal{H}$ such that $\pi(a|s,h)=1$ and $\pi(a^{\prime}|s,h)=0$ for all $a^{\prime}\in\mathcal{A}\setminus a$, we refer to $\pi$ as deterministic. Finally, for a time-dependent policy, let $\pi(s,h)$ denote the action selected deterministically while in state $s$ at time step $h$.
@@
$\pi^{*}=\mathrm{arg}\max_{\pi\in\Pi}\mathbb{E}\left[\sum\nolimits_{h=1}^{H}R(s,a,h)\right],$ (4)
is obtained by solving the recursive Bellman equations backwards-in-time according to
@@
where $V^{*}(s,h)$ is the optimal value function for state $s$ at time step $h$ and $\pi^{*}(s,h)$ is the corresponding maximizing value of $a\in\mathcal{A}$. We then evaluate the safety performance of the obtained policy $\pi$. First, we define $V^{\varphi}(s,h)=\mathrm{Pr}^{\pi}((s,h)\models\square\neg\mathcal{B}).$ Note that here $\pi$ is the policy obtained from solving the unconstrained problem (4). The safety value function $V^{\varphi}(s,h)$ is computed using the following formulation:
@@
Note that it is straightforward to generalize the constraint (13b) to the probability of multiple mission specifications. We refer the reader to Forejt et al.[9] section 8 for further details.
@@ 3 Space Environment and Dynamics
Transitioning from the abstract MDP formulation to a physically accurate representation of the operational environment is essential for synthesizing meaningful policies. While the MDP provides a framework for decision-making under uncertainty, its fidelity depends on a precise characterization of the forces governing spacecraft motion. Orbital mechanics forms the foundation for modeling a spacecraft’s trajectory by accounting for gravitational forces, atmospheric drag, and other perturbations, which in turn define the deterministic states, actions, and transitions in our MDP. In contrast, space weather introduces stochastic effects: its variability is difficult to predict yet has a direct impact on the operational environment. These uncertainties are incorporated into the transition probabilities, ensuring that the resulting policies remain robust to both predictable dynamics and unpredictable environmental fluctuations.
@@ 3.1 Orbital Mechanics
First, we introduce key ideas from orbital mechanics and then present how we can build an MDP for space mission planning. Consider a spacecraft in an elliptical orbit around the Earth. Let $\mathbf{r}(t)=[\mathbf{p}(t)^{\top},\mathbf{v}(t)^{\top}]^{\top}\in\mathbb{R}^{6}$ denote the Cartesian state of a spacecraft in the inertial frame, where $\mathbf{p}(t)=[x(t),y(t),z(t)]^{\top}$ and $\mathbf{v}(t)=[\dot{x}(t),\dot{y}(t),\dot{z}(t)]^{\top}$ are the spacecraft position and velocity at time $t$. In what follows, we will omit the time index $t$ for notational clarity. Assuming a spherical Earth, the altitude $R_{\mathrm{alt}}$ of the spacecraft above Earth’s surface can subsequently be obtained from its position $\mathbf{p}$ according to
[…]
where $R_{e}$ is the radius of the Earth. The dynamics of the spacecraft are given by
[…]
@@
where $C_{D}$ is the drag coefficient of the spacecraft, $A$ is the cross-sectional area of the spacecraft, $m$ is the mass of the spacecraft, $\rho_{A}$ is the altitude-dependent atmospheric density, $\mathbf{\textit{v}}_{a}$ is the velocity of the spacecraft relative to the Earth-fixed frame, given by
[…]
in which $\dot{\theta}$ is the rotation rate of the Earth. The atmospheric density $\rho_{A}$ is evaluated using the NRLMSISE-00 model [10, 11, 12] as a function of altitude, and the solar flux at 10.7 cm (F10.7) obtained from NOAA predictions [13].
In tandem with the preceding Cartesian representation of the spacecraft dynamics, one can instead characterize its orbit in terms of the Keplerian orbital elements $\{a,e,i,\omega,\Omega,f\}$, where $a$ is the semimajor axis length, $e\in[0,1)$ is the eccentricity, $i$ is the inclination, $\omega$ is the argument of periapse, $\Omega$ is the longitude of the ascending node, and $f$ is the true anomaly [14]. These orbital elements are visualized in Figure 2, where the angles are defined with respect to the Earth-centered inertial reference frame.
@@ 3.2 Space Weather
The solar cycle governs the Sun’s activity levels and, in turn, the atmospheric environment encountered by the spacecraft. However, the duration and intensity of individual cycles vary; larger cycles tend to last longer than 11 years, while smaller cycles are typically shorter. Solar activity is commonly categorized into three levels: high, moderate, and low [1]. During high solar activity, increased ultraviolet and X-ray radiation heat and expand the Earth’s upper atmosphere, raising atmospheric density at orbital altitudes and increasing drag forces. This accelerates orbital decay, requiring more frequent station-keeping maneuvers and increasing fuel consumption. In moderate activity periods, atmospheric density variations are less extreme, allowing for more stable orbit maintenance. Conversely, low solar activity leads to reduced drag, minimizing fuel expenditure and extending mission lifetime.